OpenStudy (anonymous):

Solve the given equation and indicate the number of values of θ sach that 0≤θ<2π satisfy the iquation cos(2θ)+cos(θ)=0

OpenStudy (vijay):

θ = 0 or π/3

OpenStudy (vijay):

cos(2θ)+cos(θ)=0 2cos^2(θ)-1+cos(θ)=0 2cos^2(θ)+cos(θ) -1=0 2cos^2(θ)-2cos(θ)+cos(θ) -1=0 2cosθ [ cosθ-1] +1 [ cos(θ-1]=0 (2cosθ-1)(cosθ-1)=0 cosθ =1/2 or cosθ = 1 θ = π/3 or θ = 0